# thanks: Corresponding author

### v1

@@ I Introduction
Electron attachment to atoms and molecules is a fundamental process in physics, chemistry, and biology, with relevance to phenomena ranging from electron transfer in photosynthesis [pshenichnyukInterconnectionsDissociativeElectron2018, rubinElectronTransportBiological1984] to radiation-induced damage in nucleic acids [narayanansjSecondaryElectronAttachmentInduced2023]. Accurate simulation of electron attachment–induced phenomena is crucial for understanding these processes holistically. The theoretical methods available for the simulation of electron affinity can be broadly classified into two distinct categories. The first category of methods consists of the so-called $\Delta$ based methods, where the electron affinity is defined as the difference between the total electronic energies of the anionic and neutral systems, both calculated at an identical level of theory. The second class of methods comprises direct energy-difference approaches, which determine the electron affinity as the transition energy between the neutral and anionic The latter methods are particularly advantageous, as they require only a single calculation, grant access to transition probabilities, and circumvent the numerical instabilities often encountered in $\Delta$-based approaches. [vooraTheoreticalApproachesTreating2017] Among the various direct energy difference-based methods available, the equation of motion coupled cluster (EOM-CC) approach is most popular due to its ssystematically improvable hierarchy. [roweEquationsofMotionMethodExtended1968, nooijenEquationMotionCoupled1995, nooijenCoupledClusterApproach1992, stantonEquationMotionCoupledcluster1993, krylovEquationofMotionCoupledClusterMethods2008]. The EOM-CC method for the electron affinity problem is generally used in singles and doubles approximation (EA-EOM-CCSD) [nooijenEquationMotionCoupled1995]. It has a formal scaling of $\mathscr{O}(\mathscr{N}^{6})$ and storage requirement which scales as $\mathscr{O}(\mathscr{N}^{4})$, where $N$ is the number of basis functions. The non-Hermitian form of the coupled-cluster similarity-transformed Hamiltonian makes property computations in the EOM-CCSD method roughly twice as costly as energy calculations. [stantonEquationMotionCoupledcluster1993]
@@
Algebraic diagrammatic construction (ADC) theory [schirmerRandomphaseApproximationNew1982, schirmerNewApproachOneparticle1983] provides a Hermitian and size-consistent alternative to the EOM-CC method, with a natural hierarchy defined by perturbation order. The ADC theory was originally developed in the context of Green’s function propagator theory. [schirmerRandomphaseApproximationNew1982] Perturbative expansions of the one-particle Green’s function—commonly referred to as the electron propagator, have given rise to a range of computational methods formulated through the Dyson equation. [vonniessenComputationalMethodsOneparticle1984, ortizElectronPropagatorPicture1997]. This group of methods is commonly referred to as the Dyson-ADC methods. Later, Schirmer et al. [schirmerNewApproachOneparticle1983] developed the non-Dyson formulation of ADC, where the electron-attachment calculation problem can be performed separately from the ionization potential calculation. The ADC equations can also be derived [banerjeeThirdorderAlgebraicDiagrammatic2019] using effective Liouvillian formalism, introduced by Mukherjee and Kutzelnigg [mukherjeeEffectiveLiouvilleanFormalism1989]. However, ADC method is generally used in Intermediate State Representation (ISR) [schirmerClosedformIntermediateRepresentations1991, mertinsAlgebraicPropagatorApproaches1996, schirmerIntermediateStateRepresentation2004, dreuwAlgebraicDiagrammaticConstruction2023]. In addition to the vertical electron-attachment energy, the $(N+1)$-electronic state wave function is also accessible in the ISR formalism. Dempwolf et al. [dempwolffIntermediateStateRepresentation2021] have reported property calculation in the EA-ADC intermediate state formalism.
@@
The second-order ADC method (EA-ADC(2)) often gives insufficient accuracy for the electron attachment problem. [schirmerNonDysonThirdorderApproximation1998, banerjeeThirdorderAlgebraicDiagrammatic2019, tripathiBoundAnionicStates2019], and the third-order approximation, EA-ADC(3), is required for higher accuracy. Similarly to the EA-EOM-CCSD method, the EA-ADC(3) method scales as $\mathscr{O}(\mathscr{N}^{6})$, but it is computationally more favorable due to its non-iterative nature of the $\mathscr{O}(\mathscr{N}^{6})$ scaling terms and lower pre-factor. [dempwolffEfficientImplementationNonDyson2019] Dreuw and his coworkers have recently explored a fourth-order approximation to the EA-ADC theory (EA-ADC(4)). [leitnerFourthOrderAlgebraicDiagrammatic2024] However, the ADC(3) method cannot be routinely used beyond systems with 10 to 15 atoms without taking additional approximations. Various strategies have been described in the literature to reduce the computational cost of wave-function-based methods. They involve approximating the two-electron integrals with the Density fitting approximation. [hattigAdvancesQuantumChemistry2005, banerjeeEfficientImplementationSinglereference2021, mukhopadhyayStatespecificFrozenNatural2023, mannaReducedCostEquation2025, mannaEfficientStateSpecificNatural2024] or use of local [pulayLocalizabilityDynamicElectron1983, pulayOrbitalinvariantFormulationSecondorder1986] and/or natural orbitals. [lowdinQuantumTheoryManyParticle1955] Among the various flavors of natural orbitals available [mesterReducedcostSecondorderAlgebraicdiagrammatic2018, duttaDomainbasedLocalPair2019, haldarEfficientFockSpace2021, duttaPairNaturalOrbital2016, landauFrozenNaturalOrbitals2010, pokhilkoExtensionFrozenNatural2020, helmichPairNaturalOrbital2013, frankPairNaturalOrbital2018, mesterReducedcostLinearresponseCC22017], the frozen natural orbital [barrNatureConfigurationInteractionMethod1970] has emerged as the most popular one. FNO-based implementation of IP [mukhopadhyayStatespecificFrozenNatural2023], DIP [mandalThirdOrderRelativisticAlgebraic2025], and EE-ADC [mesterReducedcostSecondorderAlgebraicdiagrammatic2018, mesterReducedCostSecondOrderAlgebraicDiagrammatic2023] method has been described in the literature. However, to the best of our knowledge, no natural orbital-based, low-cost ADC method is available in the literature for the electron attachment problem. The aim of this paper is to develop a low-cost ADC(3) method for the electron attachment problem based on state-specific frozen natural orbitals.
@@ II.1 Algebraic Diagrammatic Construction (ADC) Theory
In the ISR formalism [schirmerClosedformIntermediateRepresentations1991, mertinsAlgebraicPropagatorApproaches1996, schirmerIntermediateStateRepresentation2004, dreuwAlgebraicDiagrammaticConstruction2023] of ADC, a linear operator is applied on an $N$-electron correlated ground state wave function to obtain $(N+1)$-electron correlated electron-attached states
@@
where $i,j,k,$... and $a,b,c,$... denote occupied and virtual spatial orbitals, respectively.
These correlated target states are generally non-orthonormal but can be orthonormalized. First, the precursor states can be formed through Gram-Schmidt orthogonalization, which can be subsequently transformed into excitation class orthonormalized (ECO) ADC intermediate states $\left(\tilde{\psi}_{A}^{N+1}\right)$ by symmetric normalization. The ADC shifted Hamiltonian $\left(\hat{H}-E_{0}\right)$ can be expressed as a secular matrix $\left(\mathbf{M}_{AB}\right)$ in the basis of these intermediate states
@@
One can rewrite the ADC equation as an eigenvalue problem
@@
where the eigenvalues $\Omega$ are vertical electron-attachment energies and the eigenvectors $\mathbf{Y}$ act as the precursor to the spectral amplitudes. The spectral amplitudes $\mathbf{x}$ are obtained from the eigenvector $\mathbf{Y}$ as
@@
${{\mathbf{f}}_{Ap}}=\langle\tilde{\psi}_{A}^{N+1}|{{\hat{c}}_{p}}|\psi_{0}^{N}\rangle$ (7)
The secular matrix in Eq. (3) can be expanded in perturbation order, and truncation at a particular order $n$ leads to ADC($n$) equations
@@
Truncating at nn=2 leads to the ADC(2) method, nn=3 leads to the ADC(3) method. Dreuw and co-workers [bauerExploringAccuracyUsefulness2022] have proposed a semi-empirical “fractional-order” ADC scheme, where the third-order contribution to the ADC matrix ($\mathbf{C}^{(3)}$) is scaled with an empirical factor x
@@
The scaling parameter $x$ can vary from 0 to 1; the value 0 will lead to the ADC(2) method, and the value 1 will lead to the ADC(3) method. The recommended value of x is 0.5[chakrabortyRelativisticThirdorderAlgebraic2025]. The $\mathbf{M}$ matrix is generally diagonalized using Davidson’s iterative Diagonalization method [davidsonIterativeCalculationFew1975], which involves the contraction of suitably chosen trial vectors with the Hamiltonian matrix elements to construct the so-called "sigma" vectors. The programmable expressions for EA-ADC(2) and EA-ADC(3) sigma vectors are provided in the supporting information. The storage and manipulation involving two-electron integrals can be computationally expensive. One needs to use an additional approximation to reduce the storage requirements.
@@ II.2 Density Fitting (DF) Approximation
The four-centered two-electron integrals $(pq|rs)$ can be expressed as, [hohensteinDensityFittingCholesky2010]
@@
One can replace the terms ${{\phi}_{x}}(r){{\phi}_{y}}(r)$ with electron density ${\bm{\rho}}_{xy}$}
@@
The electron density ${\bm{\rho}}_{xy}$} can be fitted by an auxiliary basis as,
@@
Here, $\mathbf{d}_{P}^{xy}$denotes the fitting coefficients of the used auxiliary basis and $\mathbf{\chi}_{P}$ denotes the auxiliary basis functions. The fitting coefficients are evaluated by minimizing the functional,
@@
Therefore, the four-centered two-electron integral $(pq|rs)$ expressed in terms of three-centered two-electron integrals [dunlapApproximationsApplicationsXa2008, whittenCoulombicPotentialEnergy2003, boysAutomaticFundamentalCalculations1956] is,
@@
The three-centered two-electron integrals can be converted into a molecular orbital basis as
@@
Subsequently, the molecular-orbital integrals can be directly generated in the molecular orbital basis from the three-centered two-electron integral. In the present implementation, integrals up to two virtual indices are generated and stored, whereas integrals with three and four virtual orbitals are constructed on the fly.
@@ II.3 State-Specific Frozen Natural Orbitals (SS-FNO)
The canonical virtual orbitals are not very compact, and truncating them often leads to non-systematic error in the correlation energy. One possible solution is to transform the orbital space into the natural orbital basis, which yields systematic convergence of the correlation energy with respect to the size of the virtual space. Natural orbitals are the eigenfunctions of a correlated one-body reduced density matrix [lowdinQuantumTheoryManyParticle1955]. Among the various flavors of the natural orbitals available, [landauFrozenNaturalOrbitals2010, pokhilkoExtensionFrozenNatural2020, barrNatureConfigurationInteractionMethod1970, sosaSelectionReducedVirtual1989, taubeFrozenNaturalOrbitals2005], the frozen natural orbitals (FNO) are among the most popular. [barrNatureConfigurationInteractionMethod1970] In the FNO approximation, the occupied space is frozen at its SCF level, and the virtual space is expanded in terms of natural orbitals. The natural orbitals are obtained by diagonalizing the one-body reduced density matrix, obtained from a correlation calculation.
@@
The eigenvectors ($V$) are natural orbitals, and the corresponding eigenvalues ($\eta$) give occupation numbers of the natural orbitals. The virtual orbital with a small occupation number generally makes a very small contribution to the ground state correlation energy and can be truncated as
@@
The virtual-virtual block of the Fock matrix $\mathbf{F}$ is transformed into the truncated natural orbital basis as
@@
The matrix
@@
connects the canonical virtual molecular orbital basis to the frozen natural virtual orbital basis. One can directly transform from an atomic orbital to a frozen natural orbital basis using the transformation matrix
@@
where ${\mathbf{U}}_{occ}$} and ${\mathbf{U}}_{vir}$} are the occupied and virtual blocks of the coefficient matrix for the transformation of atomic orbitals to the canonical molecular orbitals. The second-order Møller-Plesset perturbation theory (MP2) method [mollerNoteApproximationTreatment1934] is generally used for the generation of natural orbitals. However, the reduced density generated from a correlated ground state wave function doesn’t contain enough information on the correlation in the electron-attached state, and truncation in the standard FNO basis may not yield a very systematic trend in electron affinity values. To properly describe the electron-attached state, a good first-order approximation to the electron-attached state wavefunction needs to be used for the generation of natural orbitals. Consequently, we have used the ADC(2) wave function to generate the natural orbitals. The virtual-virtual block of the one-particle reduced density matrix for the $k^{th}$ electron-attached state at the ADC(2) method has been defined as
@@
where $\mathbf{D}_{ab}^{MP2}$ and $\mathbf{D}_{ab}^{EA-ADC(2)}(k)$ denote the one-body reduced density matrices for the ground and the $k^{th}$ electron-attached state calculated at the EA-ADC(2) level of theory, respectively. The explicit expression for $\mathbf{D}_{ab}^{EA-ADC(2)}(k)$ is provided in the Supplementary Material. One can generate natural orbitals from the $\mathbf{D}_{ab}^{SS}(k)$ following the equations (20-27). The natural orbitals generated in this procedure will necessarily be state-specific. To distinguish it from the standard MP2-based FNO, it will be denoted as SS-FNO in the rest of the manuscript. It should be noted here that the one-body reduced density used for the generation of natural orbitals has been calculated using the zeroth-order intermediate state representation calculated using the EA-ADC(2) eigen vectors.

### v2

@@ I Introduction
Electron attachment to atoms and molecules is a fundamental process in physics, chemistry, and biology, with relevance to phenomena ranging from electron transfer in photosynthesis [55, 60] to radiation-induced damage in nucleic acids [47]. Accurate simulation of electron-attachment-induced phenomena is crucial for a comprehensive understanding of these processes. The theoretical methods available for the simulation of electron affinity can be broadly classified into two distinct categories. The first category of methods consists of the so-called $\Delta$-based methods, in which the electron affinity is defined as the difference between the total electronic energies of the anionic and neutral systems, both calculated at the same level of theory. The second class of methods comprises direct energy-difference approaches, which determine the electron affinity as the transition energy between the neutral and anionic systems. The latter methods are particularly advantageous, as they require only a single calculation, provide access to transition probabilities, and circumvent the numerical instabilities often encountered in $\Delta$-based approaches. [78] Among the various direct energy difference-based methods available, the equation of motion coupled cluster (EOM-CC) approach is particularly popular due to its systematically improvable hierarchy [59, 49, 50, 68, 32]. The EOM-CC method for the electron affinity problem is generally used in the singles and doubles approximation (EA-EOM-CCSD) [49]. It has a formal scaling of $\mathscr{O}(\mathscr{N}^{6})$ and storage requirements that scales as $\mathscr{O}(\mathscr{N}^{4})$, where $\mathscr{N}$ is the number of basis functions. The non-Hermitian form of the coupled-cluster similarity-transformed Hamiltonian makes property calculations in the EOM-CCSD method roughly twice as costly as the corresponding energy calculation.  [68]
@@
Algebraic diagrammatic construction (ADC) theory [64, 61] provides a Hermitian and size-consistent alternative to the EOM-CC method, with a natural hierarchy defined by perturbation order. ADC theory was originally developed in the context of Green’s function propagator theory. [64] Perturbative expansions of the one-particle Green’s function, commonly referred to as the electron propagator, have led to a variety of computational methods formulated within the framework of the Dyson equation. [74, 53]. This group of methods is commonly referred to as Dyson-ADC methods. Later, Schirmer et al. [61] developed the non-Dyson formulation of ADC, in which the electron-attachment problem can be solved separately from the ionization potential problem. The ADC equations can also be derived [1] using the effective Liouvillian formalism, introduced by Mukherjee and Kutzelnigg [45]. However, the ADC method is generally formulated in the intermediate state representation (ISR) [65, 39, 63, 14]. In addition to the vertical electron-attachment energy, the $(N+1)$-electronic state wave function is also accessible within the ISR formalism. Dempwolf et al. [11] have reported property calculations within the EA-ADC intermediate state formalism. The second-order ADC method (EA-ADC(2)) often provides insufficient accuracy for the electron attachment problem [62, 1, 73], and the third-order approximation, EA-ADC(3), is required for higher accuracy. Similar to the EA-EOM-CCSD method, the EA-ADC(3) method scales as $\mathscr{O}(\mathscr{N}^{6})$, ), but it is computationally more favorable due to the non-iterative nature of the $\mathscr{O}(\mathscr{N}^{6})$ scaling terms and its lower pre-factor. [12] Dreuw and his coworkers have recently explored a fourth-order approximation to the EA-ADC theory, denoted as EA-ADC(4). [34]
@@
The ADC(3) method cannot be routinely applied to systems containing more than 10–15 atoms without additional approximations. Various strategies have been described in the literature to reduce the computational cost of wave-function-based methods. These strategies involve approximating the two-electron integrals using the density-fitting approximation  [24, 2, 46, 37, 38] or employing local  [57, 56] and/or natural orbitals. [35] Among the various flavors of natural orbitals available [43, 18, 22, 17, 33, 54, 26, 19, 42], the frozen natural orbital [4] have emerged as the most popular approaches. FNO-based implementation of IP [46], DIP [36], and EE-ADC [43, 40] method have been described in the literature. However, to the best of our knowledge, no natural-orbital-based low-cost ADC method has been reported for the electron-attachment problem. The aim of this work is to develop a low-cost ADC(3) method for the electron attachment problem based on state-specific frozen natural orbitals.
@@ II.1 Algebraic Diagrammatic Construction (ADC) Theory
In the ISR formalism of ADC,  [65, 39, 63, 14] electron-attached states are generated by applying a linear operator to the correlated $N$-electron ground-state wave function.
@@
where $i,j,k,$...denote occupied orbitals and $a,b,c,$... denote virtual orbitals.
These correlated target states are generally non-orthonormal and must therefore be orthonormalized. First, precursor states are constructed through Gram–Schmidt orthogonalization. These states are then transformed into excitation-class-orthonormalized (ECO) ADC intermediate states by symmetric normalization. The ADC shifted Hamiltonian $\left(\hat{H}-E_{0}\right)$ can be expressed as a secular matrix $\left(\mathbf{M}_{AB}\right)$ in the basis of these intermediate states
@@
The ADC equation can be written as the eigenvalue problem
@@
where the eigenvalues $\Omega$ are vertical electron-attachment energies and the eigenvectors $\mathbf{Y}$ determine the spectral amplitudes. The spectral amplitudes $\mathbf{x}$ are obtained from the eigenvector $\mathbf{Y}$ as
@@
${{\mathbf{f}}_{Ap}}=\langle\tilde{\psi}_{A}^{N+1}|{{\hat{c}}_{p}^{\dagger}}|\psi_{0}^{N}\rangle$ (7)
The secular matrix in Eq. (3) can be expanded in perturbation order, and truncation at order $n$ leads to ADC($n$) equations
@@
Truncating at nn=2 yields ADC(2) method, whereas truncation at nn=3 yields ADC(3). Dreuw and co-workers [6] have proposed a scaled matrix (sm) “fractional-order” ADC scheme, where the third-order contribution to the ADC matrix is scaled with an empirical factor x
@@
The scaling parameter $x$ can vary from 0 to 1; the value 0 will lead to the ADC(2) method, and the value 1 will lead to the ADC(3) method. The recommended value of x is 0.5[9]. The $\mathbf{M}$ matrix is generally diagonalized using Davidson’s iterative diagonalization method [10], which involves the contraction of suitably chosen trial vectors with the Hamiltonian matrix elements to construct the so-called "sigma" vectors. The programmable expressions for EA-ADC(2) and EA-ADC(3) sigma vectors are provided in the Supplementary Material. The storage and manipulation of terms involving two-electron integrals can be computationally expensive and requires large storage, especially in extended basis sets. One needs to use an additional approximation to reduce the storage requirements.
@@ II.2 Density Fitting (DF) Approximation
The four-centered two-electron integrals $(pq|rs)$ are defined as  [27]
@@
The product of orbital functions ${{\phi}_{x}}(r){{\phi}_{y}}(r)$ can be interpreted as an orbital pair density ${\bm{\rho}}_{xy}$}
@@
The pair density ${\bm{\rho}}_{xy}$} can be expanded in an auxiliary basis as
@@
Here, $\mathbf{d}_{P}^{xy}$denotes the fitting coefficients and $\mathbf{\chi}_{P}$ denotes the auxiliary basis functions. The fitting coefficients are determined by minimizing the Coulomb metric functional,
@@
Therefore, the four-centered two-electron integral $(pq|rs)$ can be expressed in terms of three-centered integrals as  [15, 80, 8],
@@
The three-centered integrals can be converted into a molecular orbital basis as
@@
Subsequently, molecular orbital integrals can be constructed directly from the three-center integrals. In the present implementation, integrals up to two virtual indices are generated and stored, whereas integrals with three and four virtual orbitals are constructed on the fly.
@@ II.3 State-Specific Frozen Natural Orbitals (SS-FNO)
Canonical virtual orbitals are generally not compact, and truncating them often leads to non-systematic errors in the correlation energy. A common approach is to transform the virtual orbital space into a natural orbital basis, which provides systematic convergence of the correlation energy with respect to the size of the virtual space. Natural orbitals are defined as the eigenfunctions of a correlated one-particle reduced density matrix [35]. Among the various types of natural orbitals, [33, 54, 4, 67, 72], the frozen natural orbitals (FNO) are the most popular. [4] In the FNO approximation, the occupied orbitals are kept frozen at the SCF level, while the virtual space is transformed to the natural orbital basis. Natural orbitals are obtained by diagonalizing the one-particle reduced density matrix derived from a correlated calculation.
@@
where $D$ is the virtual-virtual block of the one-particle density matrix. The eigenvectors ($V$) are natural orbitals, and the corresponding eigenvalues ($\eta$) give occupation numbers of the natural orbitals. The virtual orbital with a small occupation number generally makes a very small contribution to the ground state correlation energy and the natural orbital with occupation below a critical threshold ($\eta_{crit}$) can be truncated as
@@
The virtual-virtual block of the Fock matrix $\mathbf{F}$ is transformed to the truncated natural orbital basis as
@@
to obtain semi-canonical orbitals. The matrix
@@
connects the canonical virtual molecular orbital basis to the frozen natural virtual orbital basis. The transformation from the atomic orbital basis to the frozen natural orbital basis can be written as
@@
where ${\mathbf{U}}_{occ}$} and ${\mathbf{U}}_{vir}$} are the occupied and virtual blocks of the coefficient matrix for the transformation of atomic orbitals to the canonical molecular orbitals. The second-order Møller-Plesset perturbation theory (MP2) method [44] is generally used for the generation of natural orbitals. However, the reduced density obtained from the correlated ground-state wave function does not contain sufficient information about correlation in the electron-attached state. Consequently, truncation in the standard FNO basis may not yield systematic convergence of electron affinity values. To properly describe the electron-attached state, a first-order approximation to the electron-attached wave function should be used to generate natural orbitals. Consequently, we have used the EA-ADC(2) wave function to generate the natural orbitals. The virtual-virtual block of the one-particle reduced density matrix for the $k^{\text{th}}$ electron-attached state in the ADC(2) method has been defined as
@@
where $\mathbf{D}_{ab}^{MP2}$ and $\mathbf{D}_{ab}^{EA-ADC(2)}(k)$ denote the one-particle reduced density matrices corresponding to the ground and the $k^{\text{th}}$ electron-attached state calculated at the MP2 and EA-ADC(2) level of theory, respectively. The explicit expression for $\mathbf{D}_{ab}^{EA-ADC(2)}(k)$ is provided in the Supplementary Material. Natural orbitals are generated from the $\mathbf{D}_{ab}^{SS}(k)$ following the equations (20-27). The natural orbitals generated in this procedure will necessarily be state-specific. To distinguish it from the standard MP2-based FNO, it will be denoted as SS-FNO in the rest of the manuscript. It should be noted that the one-particle reduced density matrix used for generating natural orbitals is constructed using the zeroth-order intermediate-state representation derived from the EA-ADC(2) eigenvectors.